# SemiHMER: Semi-supervised Handwritten Mathematical Expression Recognition using pseudo-labels

### v1

@@ 4.5.2 Strategy of the Weak-to-Strong augmentation
To allow the model to benefit from both data augmentation and cross-pseudo supervision strategies, we conducted experiments with different augmentation intensities. A natural idea is to gradually increase the intensity of augmentation. As shown in Table 3, by progressively introducing warping, stretching, and perspective transformations into the model, performance on the dataset CROHME2014 improved by 0.09%, 0.58%, and 2.02% respectively. Finally, when applying both weak-strong augmentations and cross-pseudo supervision strategies to both branches, the results further improved by 1.92%, 2.13%, and 2.63%, demonstrating the effectiveness of our strategy: as the augmentation becomes more effective, the impact of semi-supervision becomes more pronounced.
@@
We ablated each component of SemiHMER step by step. Table 4 shows that applying weakly augmented samples to both branches directly results in a 1.53% decrease in performance. This is possible because both branches only introduce unhelpful noisy errors to the model when without augmentation. Naturally, we considered that simultaneously applying strong augmentation to both branches, enhancing their respective capabilities, would improve the credibility of the supervision signal.
However, the result was only a 1.3% gain, which might be attributed to the fact that simultaneous strong augmentation would exponentially increase the amount of erroneous supervisory information introduced by both branches. Finally, our method, which uses a strategy of weak augmentation on one branch and strong augmentation on the other branch, improves performance by 4.65%, demonstrating that the proposed weak-to-strong augmentation strategy for cross-pseudo supervision is a more powerful tool for semi-supervised handwritten mathematical expression recognition.
4.5.3 The trade-off weight $\lambda$
In the experiments above, the symbol $\lambda$ is used to balance the trade-off between supervised loss and unsupervised loss. Table 5 shows that $\lambda=10^{-3}$ performs best in our setting, where a smaller $\lambda=10^{-4}$ reduces much of the useful information provided by the pseudo labels. A larger $\lambda=10^{-2}$ is problematic and leads to performance degradation because the network might converge in the wrong direction.
Table 5: Performance analysis of different weight $\lambda$
Value of $\lambda$ $\lambda=10^{-2}$ $\lambda=10^{-3}$ $\lambda=10^{-4}$
@@ 5 Conclusion
While handwritten mathematical expression recognition has been well developed over the past few decades, the effectiveness of consistency regularization and self-training has not received much attention. Therefore, how to exploit the potential benefits of noisy pseudo labels while utilizing semi-supervised learning is of great significance for future research.

### v3

@@ 4.7 Effectiveness of the weak-to-strong augmentation strategy
We ablated each component of SemiHMER step by step. 4, shows that directly applying weakly augmented samples to both branches results in a 1.53% decrease in performance, which is possible because, without augmentation, both branches only introduce unhelpful noisy errors to the model. Naturally, we considered that simultaneously applying strong augmentation to both branches, enhancing their respective capabilities, would improve the credibility of the supervision signal. However, the result was only a 1.3% gain, which might be attributed to the fact that simultaneous strong augmentation would exponentially increase the amount of erroneous supervisory information introduced by both branches. Finally, our method, which uses a strategy of weak augmentation on one branch and strong augmentation on the other branch, improved the model’s performance by 4.65%, demonstrating that the proposed weak-to-strong augmentation strategy for cross-pseudo supervision is a more powerful tool for semi-supervised handwritten mathematical expression recognition.
@@
4.8 The trade-off weight $\lambda$
$\lambda$ is used to balance the trade-off between supervised loss and unsupervised loss. Table 5 shows that $\lambda=10^{-3}$ performs best in our setting, where a smaller $\lambda=10^{-4}$ reduces much of the useful information provided by pseudo-labels. A larger $\lambda=10^{-2}$ is problematic and leads to performance degradation because the network could converge in the wrong direction.
Table 5: Performance analysis of different weight $\lambda$.
$\lambda$ $10^{-2}$ $10^{-3}$ $10^{-4}$
@@ 5 Conclusion
Although handwritten mathematical expression recognition has been well developed over the past few decades, the effectiveness of consistency regularization and self-training has not received much attention. Therefore, how to exploit the potential benefits of noisy pseudo-labels while utilizing semi-supervised learning is of great significance for future research.